The average energy of particles in the early phase of the universe, when it was very
hot, was approximately 10 GeV. Calculate the corresponding temperature.
1
Expert's answer
2019-04-29T09:29:38-0400
"10 \ GeV = 1.6 \cdot 10^{-9} J"

We need to use the formula for the energy:


"E=3kT"

"T=\frac{E}{3k}=\frac{1.6 \cdot 10^{-9} }{3(1.38 \cdot 10^{-23})}=3.9 \cdot 10^{13} \ K"